Please help.If w= sm-8/6m, w=1.4, and s=10, find m.
Answer by harpazo(655)

Please help.If w= sm-8/6m, w=1.4, and s=10, find m.
1.4 = (10m - 8)/6m
1.4 * 6m = 10m - 8
8.4m = 10m - 8
8.4 - 10m = -8
-1.6m = -8
m = -8/-1.6
m = 5
Done!